OpenStudy (anonymous):

Divide 5x^2y-8x^4y^2+9y^3 / 2xy^2

OpenStudy (gw2011):

If you divide each by 2xy^2 you get: 5x^2y/2xy^2=5x/2y -(8x^4y^2/2xy^2)=-4x^3 9y^3/2xy^2=9y/2x Combining the above you get: 5x/2y-4x^3+9y/2x

OpenStudy (anonymous):

factor out a y in the top fraction and then reduce the y's

OpenStudy (anonymous):

so how do i put that in the computer?

OpenStudy (anonymous):

can u show me on white board how to enter that?